OpenStudy (anonymous):

This thing... 1+x+x^2+x^3+...=1/(1-x) Is true for |x| < 1 right?

OpenStudy (anonymous):

If so why is Euler (in step 8) allowed to do that? http://math.ucr.edu/home/baez/qg-winter2004/zeta.pdf

OpenStudy (akashdeepdeb):

This is known as an infinite series. Hold on i'll check the link.

OpenStudy (akashdeepdeb):

http://www.youtube.com/watch?v=w-I6XTVZXww

OpenStudy (anonymous):

Yeah that's where I got it from... there's a linked vid where he does a better proof. In it he differentiates the above, then plugs in x = -1, which isn't in the domain...

OpenStudy (akashdeepdeb):

Not sure then.

OpenStudy (anonymous):

Haha, thanks anyway!

OpenStudy (anonymous):

@ParthKohli any thoughts?

Parth (parthkohli):

You are right about that. Even though I read the document, I'm sorry to say that I have not reached that level of mathematics.

OpenStudy (akashdeepdeb):

much truth. very amaze wow

Parth (parthkohli):

The document accepts that the series diverges, but it is using the equation from step 7.

OpenStudy (akashdeepdeb):

doge much smart

OpenStudy (anonymous):

Here's that video I'm talking about http://www.youtube.com/watch?v=E-d9mgo8FGk In it he says it's true for x<1

OpenStudy (anonymous):

is that chinese

OpenStudy (anonymous):

i know it is math but what kind of math is THAT!!!

Parth (parthkohli):

I've watched only one proof. Didn't know there was a second one too!

OpenStudy (amoodarya):

\[s(n) =1+x+x^2+x^3+...+x ^{n-1}\] \[x \times s(n) =1+x+x^2+x^3+...+x ^{n-1} =x+x^2+x^3+...+x ^{n-1}+x ^{n1}\] now s(n) -x s(n) =\[x-s(n) =1 -x ^{n} \rightarrow s(n)=\frac{1 -x ^{n} }{1-x }\] \[\lim_{n \rightarrow \infty} s(n)=\lim_{n \rightarrow \infty}\frac{1 -x ^{n} }{1-x }=\frac{1 }{1-x } \rightarrow \lim_{n \rightarrow \infty}x^n=0 \rightarrow |x|<1 \]

OpenStudy (anonymous):

"By using the equation in part 7 to evaluate the formal power series" formal power series exist independently of considerations of convergence

OpenStudy (anonymous):

anyways the Abel sum is simple:$$A\sum_{n=1}^\infty (-1)^{n+1} n= \lim_{t\to 1^-}\sum_{n=1}^\infty (-1)^{n+1}nt^n=-\lim_{t\to 1^-}\sum_{n=1}^\infty n(-t)^n$$

OpenStudy (anonymous):

now we know:$$\sum_{n=1}^\infty x^n=\frac{x}{1-x},|x|<1\\\frac{d}{dx}\sum_{n=1}^\infty x^n=\frac{d}{dx}\frac{x}{1-x}\\\sum_{n=1}^\infty\frac{d}{dx}x^n=\frac1{(1-x)^2}\\\sum_{n=1}^\infty nx^{n-1}=\frac1{(1-x)^2}\\\sum_{n=0}^\infty(n+1)x^n=\frac1{(1-x)^2}\\\sum_{n=0}^\infty nx^n+\sum_{n=0}^\infty x^n=\frac1{(1-x)^2}\\\sum_{n=0}^\infty nx^n=\frac1{(1-x)^2}-\frac1{1-x}$$ergo we find:$$-\lim_{t\to1^-}\left(\frac1{(1+t)^2}-\frac1{1+t}\right)=\frac14$$

OpenStudy (anonymous):

thus \(\displaystyle A\sum_{n=1}^\infty(-1)^{n+1}n=\frac14\)

OpenStudy (anonymous):

note in the sum we used \(x\mapsto -t\)